Find the solution of the given differential equation and then find the particular solution for which a point (x,y) is given:


dy/dx = (3x - 4)


1
Expert's answer
2021-08-30T12:41:12-0400

"\frac{dy}{dx}=3x-4\\
dy=(3x-4)dx\\
\text{Integrating both side, we get}\\
y=\frac{3}{2}x^2-4x+c\\
\text{This is the required general solution.}\\
\text{To find particular solution we need one initial condition.}"